How should I calculate the angle by which two wheels are rotated driving the vehicle around the circle. I do know circle radius, wheel radius, car length and distance between wheels. Thanks


1
Expert's answer
2021-01-01T14:07:32-0500




"\text{the car moves in a circle, therefore all points of the car move in}"

"\text{circles with one common center}"

"\text{let } r -\text {wheel radius}"

"R -\text{radius of the circle of movement}"

"l - \text{car length}"

"w -\text{car width}"

"\text{also assume that the rear wheels of the car}"

"\text{do not turn}"

"\alpha_1,\alpha_2 \text{ angles of inclination of the front wheels}"

"L - \text{distance between axles car}"

"L= l-2r"

"\sin(\alpha_1)= \frac{L}{R}=\frac{l-2r}{R}"

"\sin(\alpha_2)= \frac{L}{R-w}=\frac{l-2r}{R-w}"

"\text{we assume that the radius of the circle of movement }"

"\text{is much greater than the width of the car}"

Answer: "sin(\alpha)=\frac{l-2r}{R}"